# MetaphorShare: A Dynamic Collaborative Repository of Open Metaphor Datasets

### v1

@@ Abstract
The metaphor studies community has developed numerous valuable labelled corpora in various languages over the years. Many of these resources are not only unknown to the NLP community, but are also often not easily shared among the researchers. Both in human sciences and in NLP, researchers could benefit from a centralised database of labelled resources, easily accessible and unified under an identical format. To facilitate this, we present MetaphorShare, a website to integrate metaphor datasets making them open and accessible. With this effort, our aim is to encourage researchers to share and upload more datasets in any language in order to facilitate metaphor studies and the development of future metaphor processing NLP systems. The website is accessible at www.metaphorshare.com.
@@ 1 Introduction
Figure 1: MetaphorShare search page.
@@
On one side, numerous dataset labelled with metaphoric usage of words are created by researchers in many languages every year, and remain out of the scope of NLP research and evaluations of the systems because they are not publicly shared or consistently formatted. On the other, metaphor studies researchers often work on domain specific corpora, languages other than English (for which NLP has developed the most resources), and use different definitions or retain different types of metaphors for different projects, with no convenient way to systematically compare their labels with other researchers, or to rely on existing metaphor an automatic labelling tool that answer a project specific need. Moreover, these created datasets are not used by other communities or are often labeled using different schemata and formats.
We propose to help speeding up collaborations between AI/NLP communities and linguistics/metaphor studies by unifying dataset formats and facilitating access to existing resources by everyone. MetaphorShare is a website designed at sharing new and publicly available existing resources (see Figure 1). It is an online dynamic repository where anyone can upload or download open metaphor datasets. Our hope is that the website will create a new synergy between different components of metaphor studies, limit the loss of datasets, catalog them and facilitate their automatic processing through a unified format. The targeted audience is anyone running corpus based research on metaphors, including NLP researchers for the development of metaphor processing and identification systems. [^1] An introduction to MetaphorShare is available at https://youtu.be/e8IE7RJPA0Q
Dataset Reference License N N dist. N dist. % Expr. Domain/Source
Name ctxt. expr. met. PoS
Psycholinguistics
JANK Jankowiak (2020) CC BY 4.0 240222Anomalous sentences are not shown in the table. 240 120 50 N constructed examples
CARD_V Cardillo et al. (2010) CC BY-NC 280 280 140 50 V constructed examples
CARD_N Cardillo et al. (2010, 2017) CC BY-NC 512 512 256 50 N constructed examples
NLP
MOH Mohammad et al. (2016) see data page [^3] https://saifmohammad.com/WebPages/metaphor.html 1632444the original dataset contains 1639 instances. A few duplicated example sentences caused by orthographic variants of the target word, such as distil/distill have been removed. 1632 439 25 V WordNet examples
NewsMet Joseph et al. (2023) Apache-2.0 1205555We show the manually labelled sentences (named gold by the authors) 1205 477 49 V Fake News Corpus [^6] https://github.com/several27/FakeNewsCorpus
TSV_A Tsvetkov et al. (2014) see data page [^7] https://github.com/ytsvetko/metaphor/blob/master/LICENSE.md 1945 1072 687 50 A various websites
GUT Gutiérrez et al. (2016) AFL-3.0 8591 3479 23 54 A Wikipedia, UKWaC [^8] Ferraresi et al. (2008)
Multi-Word Expressions
PVC Tu and Roth (2012) no license [^9] https://cogcomp.seas.upenn.edu/page/resource_view/26 1348 1348 23 65 V-Prep BNC
MAD Tayyar Madabushi et al. (2022b) GPL-3.0 4558 4554 251 48 NC Common Crawl
MAGPIE Haagsma et al. (2020) $C$-BY-4.0 48395 47283 9307 75 Various BNC, PMB [^10] Abzianidze et al. (2017).
MIPVU
VUAC_BO Boisson et al. (2023) CC BY-SA 3.0 39223 11476 8674 52 Various VUAC
TONG Tong et al. (2024) $C$-BY-4.0 1428111111Original VUAC sentences and apt (original dataset label) paraphrases are counted in the table. 739 861 46 Various VUAC & paraphrases
Table 1: Twelve example metaphor datasets included in the MetaphorShare repository. The N column shows the number of instances in the datasets, followed by the number of distinct provided context surrounding the potential metaphoric expression (N dist ctxt), and the number of distinct lemmas labelled metaphoric or literal (N dist. expr.). %met. indicates the percentage of expressions labelled as metaphorical, and Expr. PoS their part of speech.
@@ 2 Related work
Veale and Li (2012) propose a system of metaphor interpretation and generation, Metaphor Magnet, that relies on the harvest of stereotype from the Google n-grams. The stereotypical attributes associated to concepts are leveraged for the suggestion of relevant metaphors given input entities such as Google is –Microsoft, and for the interpretation of the metaphor. More recently, Mao et al. (2023) release end-to-end metaphor identification and interpretation website for free text (non-domain specific): MetaPro tags metaphoric tokens or multi-word expressions in sentences, generates paraphrases, and outputs a more abstract concept mapping derived from the analysis of the sentence. The authors list limitations due to the lack of training data for metaphor inference and extended metaphors. The main difference with these two initiatives and ours is that our aim is to integrate already labeled datasets (by experts or others) into a unified repository, encouraging re-sharing of resources, rather to provide a platform to annotate new datasets automatically. Similarly to ours, MetaPro also uses several metaphor datasets for training the models Mao et al. (2019), but in this case they do not mention more than six from the NLP community.
@@ 3 Data Collection and Unification
MetaphorShare aims at facilitating computational research on multiple datasets. A minimally constrained format is required to store the datasets into a database, conveniently compare the resources, search efficiently into the different fields across datasets, and run experiments on several sets at once. Metaphor datasets can be added by their authors or by anyone if their license allows it. In this section, we review the existing dataset formats and our unification choices.
We select ten open metaphor datasets with open licenses and diverse encoded information to illustrate our unification process and as representative examples for our repository. Their characteristics are summarized in Table 1.
One expression per example.
@@
Many datasets in NLP contain labels for binary classification or paraphrasing of one delimited expression within a given context. The expressions considered might be a single token, a multi-word expression (MWE), a phrase or a compound. The task of their original release paper may be to decide if the marked expression is used metaphorically or literally, or to find a correct literal paraphrase of the metaphoric expression, among other possible analysis and tasks. The context may consist in one single word, in the case of Adjective-Noun pairs datasets, as for example in Gutiérrez et al. (2016) or Tsvetkov et al. (2014). They often consist in one sentence such as Mohammad et al. (2016) or Cardillo et al. (2010, 2017) or Jankowiak (2020). Longer contexts are sometimes directly provided, such as in Tayyar Madabushi et al. (2022b) or Haagsma et al. (2020) who include two sentences before and after the sentence containing the marked Potential Metaphoric Expression (PME), or indirectly provided with a reference to the original document, as for example in Joseph et al. (2023) who share references to newspaper articles or Tu and Roth (2012) who link the annotated examples to sentence pointers in the British National Corpus.
The Metaphor Identification Procedure (MIPVU).
@@
Crisp et al. (2007) and then Steen (2010) defined an annotation guideline that is widely used in metaphor studies, with the release of the VU Amsterdam Corpus (VUAC), the largest metaphor corpus built in recent years. A few other examples of English open datasets created following MIPVU are Nacey et al. (2019b) and Nina Julich-Warpakowski’s work on the Music Criticism domain [^12] https://osf.io/scfqu/. Many other MIPVU datasets exist in various languages with adaptations of the guideline Nacey et al. (2019a), for example Nacey (2022) for Scandinavian.
@@
In this framework, every token in a document is labelled as metaphoric or literal, with more fine grain characterization of the metaphoric tokens. Conventional metaphors, such as prepositions (e.g. in my mind) are also considered metaphoric. Different versions of the VUAC suiting NLP tasks, adapted for binary classification tasks Leong et al. (2018), or including additional ratings of novelty Parde and Nielsen (2018) have been released over the years. In the evaluation section of this paper, Section 5, we use a modified version of the VUAC corpus, adapted for binary classification (VUAC_BO) as described in Boisson et al. (2023), where one instance contains information about one labelled expression. The same sentence may be duplicated with different PME of focus.
@@
Beyond metaphorical/literal labels of expressions in context, metaphor datasets may contain more or less common additional information. There are scalar variables such as metaphoricity/figurativeness ratings Cardillo et al. (2010); Dunn (2014); Katz et al. (1988), confidence scores, concreteness scores, novelty scores Parde and Nielsen (2018), level of emotion scores Mohammad et al. (2016), frequency in corpora… They may also contain categorical variables such as PoS, textual source and genre, metaphor type, source and target concept/domains information Gordon et al. (2015). Figure 2 in the appendix shows an example of the additional information encoded in the Mohammad et al. (2016) dataset for one instance, as diplayed on MetaphorShare. Some datasets, for example the LCC Mohler et al. (2016) and Gordon et al. (2015), also annotate the fragment of the context that carries the lexical information about the target of a metaphor (e.g. in the phrase ocean of happiness, the metaphoric expression is ocean and the target lexical cue of the metaphor is happiness).
@@
To simplify the search into various datasets and their handling for further system development, we require a specific input format for uploading a new resource on the website. MetaphorShare currently accepts CSV files having minimally one column named tagged_sentence. The column must contain text with at least one tagged expression per line. The accepted tag set is currently: < m: metaphoric, l: literal, t: target, u: free tag >. Each cell must minimally contain one tag and multiple tags are allowed, making possible the upload of MIPVU datasets without sentence duplication, e.g. I <m>swim</m> today in an <m>ocean</m> of <t>happiness</t>. Additional information may be added in freely named columns of the CSV file. Sentence indices can therefore be added to preserve the sentence order in a document and include metaphor in discourse datasets. Redundant fields across datasets such as PoS information have recommended names provided in the website guideline and we will see in Section 4.2 that the manual validation step by the website administrators allows a discussion with the user/author for further field names unification if needed. If any author of a dataset encounters issues to reformat a dataset into our expected format, we invite her or him to contact us by email for support.
@@
Author(s) Authors of the paper/dataset
@@
Source Source docu of the labelled instances
@@
Comments Any additional description and comments
@@ 4 MetaphorShare: The Website
MetaphorShare general architecture is described in Section 4.1. The website is organized around three main pages: an uploading page (Section 4.2), a dataset catalog (Section 4.3) and a page for searching into the datasets records (Section 4.4).
@@ 4.2 Uploading a new dataset
The required and optional fields to be filled in at uploading time are presented in Table 2. For example, if a dataset is associated with one or several publications, the references should be added in the appropriate fields, preferably in the BibTeX format. The information provided will be displayed in the dataset catalog, and will help to further categorize the datasets in the future versions of the website.
@@
After validating the form, an immediate automatic file format check on the data is done and feedback is returned to the user for acceptance or rejection of the file. Outputs of this step can be seen in Appendix Figure 4 & 5. In a second manual validation step, the website administrators check the license information, dataset reference, and eventually suggest modifications of CSV free columns names, that contain the dataset specific additional information, in an email interaction with the user, for database field unification purposes.
@@
Once a dataset passes the two validation steps, it is stored in the main repository database. This ElasticSearch database has two indices : the Datasets index similar to Table 2 fields stores all the user provided metadata, and information extracted during the processing of the data file such as the dataset size and the label distribution. The second one, the PME index, contains the output of the data file parse, as shown Table 3. Each textual labelled example is linked to the Dataset index.
@@ 4.3 Dataset Catalog
A partial view of a dataset summary in the catalog is shown in the Appendix, Figure 3. The fields filled at uploading time, including the license chosen, are displayed. Statistics computed from the input file are also shown and plotted, in particular the number of instances, label set and distribution. The list of additional fields is also extracted from the input file.
@@ 4.4 Search Page
The records can be accessed and downloaded through a search page, shown in Figure 1. The search fields include optional selections of datasets and languages from dropdown lists, restricted search in a label (metaphorical or literal), and keyword search into the PME text (with exact and fuzzy match), or into the entire text. One phrase or sentence is shown by default on the result page, but longer context can be accessed in one click, as well as all the additional information attached to a record. The organization of the catalog and the fine-grain search options are likely to evolve in future versions of MetaphorShare.
@@ 5 Evaluation
As a case study showing the potential research that can be carried out using our centralised repository, we perform a cross-dataset metaphor identification analysis. We evaluate the datasets presented Table 2, as downloaded from the repository. The ten unified datasets are then used in a cross-dataset binary classification experiment to illustrate a possible usage of MetaphorShare for NLP research. Given an input expression, the task consists in deciding whether the expression is a metaphor or not. Different projects label metaphors for different purposes. MetaphorShare makes it easy to fine-tune models on relevant dataset to further support project-specific automatic or semi automatic labeling.
@@
In order to maintain a training set of the same size for all datasets, we randomly sample 800 examples from each set to create the training sets. Because our datasets from psycholinguisics are small and similar in their creation method, we group them into one single set called J&C in this experiment. RoBERTa base models Liu et al. (2019) are then finetuned independently on the 10 datasets and tested on each of them. Similarly to the experiments in Boisson et al. (2023), for hyperparameter optimisation, we rely on the Bayesian Optimization with Hyperband (BOHB) algorithm Falkner et al. (2018) available in RayTune Liaw et al. (2018), with an identical hyperparameter search space [^13] The hyperparameters search space is set to a batch size equal to 4, 8 or 16; the learning rate in a ranging from 5e-7 to 5e-5; the number of epochs within 1 to 12; and the random seeds taking three possible values (1, 2 and 3). and 25 trials.

### v2

@@ Abstract
The metaphor studies community has developed numerous valuable labelled corpora in various languages over the years. Many of these resources are not only unknown to the NLP community, but are also often not easily shared among the researchers. Both in human sciences and in NLP, researchers could benefit from a centralised database of labelled resources, easily accessible and unified under an identical format. To facilitate this, we present MetaphorShare, a website to integrate metaphor datasets making them open and accessible. With this effort, our aim is to encourage researchers to share and upload more datasets in any language in order to facilitate metaphor studies and the development of future metaphor processing NLP systems. The website has four main functionalities: upload, download, search and label metaphor datasets. It is accessible at www.metaphorshare.com.
@@ 1 Introduction
Figure 1: MetaphorShare search page. Specific datasets, languages, and tag types can be selected, and a text-based search within tagged expressions or into the entire text is implemented. Additional features provided with the record appear when clicking the Show Details button.
@@
On one side, numerous datasets labelled with metaphoric usage of words are created by researchers in many languages every year, and remain out of the scope of NLP research and evaluations of the systems because they are not publicly shared or consistently formatted. On the other, metaphor studies researchers often work on domain specific corpora, languages other than English (for which NLP has developed the most resources), and use different definitions or retain different types of metaphors for different projects, with no convenient way to systematically compare their labels with other researchers, or to rely on existing metaphor an automatic labelling tool that answer a project specific need. Moreover, these created datasets are not used by other communities or are often labeled using different schemata and formats.
We propose to help speeding up collaborations between AI/NLP communities and linguistics/metaphor studies by facilitating the unification of dataset formats and the access to existing resources by everyone, while preserving the information encoded in the original datasets. MetaphorShare is a website designed at sharing new and publicly available existing resources (see Figure 1). It is an online dynamic repository where anyone can upload or download open metaphor datasets. Our online labeling tool also offers the possibility to directly create a dataset, and optionally share it. Our hope is that the website will create a new synergy between different components of metaphor studies, limit the loss of datasets, catalog them and facilitate their automatic processing through a unified format. The targeted audience is anyone running corpus based research on metaphors, including NLP researchers for the development of metaphor processing and identification systems. [^1] An introduction to MetaphorShare is available at https://youtu.be/Fi48SOjueEE
@@ 2 Related work
Veale and Li (2012) develop a system of metaphor interpretation and generation, Metaphor Magnet, that relies on the harvest of stereotype from the Google n-grams. The stereotypical attributes associated to concepts are leveraged for the suggestion of relevant metaphors and their interpretation (e.g. given Google is –Microsoft, the system outputs giant with properties like lumbering and sprawling). More recently, Mao et al. (2023) release end-to-end metaphor identification and interpretation website for free text (non-domain specific): MetaPro tags metaphoric tokens or multi-word expressions in sentences, generates paraphrases, and outputs a more abstract concept mapping derived from the analysis of the sentence. The authors list limitations due to the lack of training data for inference and extended metaphors. The main difference with these two initiatives and ours is that our aim is to integrate already labeled datasets into a unified repository, encouraging the structuring of resources and expression of needs from the metaphor study community, before building automatic annotation tools. Similarly to ours, MetaPro also uses several metaphor datasets for training the models Mao et al. (2019), but in this case they do not mention more than six from the NLP community.
Annotation tools.
Many open source (e.g. Label Studio, Dataturks, Doccano Nakayama et al. (2018) and Potato Pei et al. (2022)) and proprietary (e.g. LabelBox, Prodigy or Amazon Mechanical Turk) annotation tools exist. Our purpose is not to compete with these services but rather to offer a simple targeted experience for metaphor annotation, in which the output is directly mapped to the unified repository format. This way, non-expert users can avoid the need for learning a general tool, as well as having to learn how to modify given labeled outputs, which wouldn’t be trivial for most.
@@
Name Reference N %M
Psycholinguistics
JANK Jankowiak (2020) 24022footnotemark: 2 50
CARD_V Cardillo et al. (2010) 280 50
CARD_N Cardillo et al. (2010, 2017) 512 50
NLP
MOH Mohammad et al. (2016) 1632 25
NewsMet Joseph et al. (2023) 1205 49
TSV_A Tsvetkov et al. (2014) 1945 50
GUT Gutiérrez et al. (2016) 8591 54
Multi-Word Expressions
PVC Tu and Roth (2012) 1348 65
MAD Tayyar Madabushi et al. (2022b) 4558 48
MAGPIE Haagsma et al. (2020) 48395 75
MIPVU
VUAC_BO Boisson et al. (2023) 39223 52
TONG Tong et al. (2024) 1428 46
Table 1: Twelve metaphor datasets included in the MetaphorShare repository. The N column shows the number of instances in the datasets. %M indicates the percentage of expressions labelled as metaphorical.
@@ 3 Data Collection and Unification
MetaphorShare aims at facilitating computational research on multiple datasets. A minimally constrained format is required to store the datasets into a database, conveniently compare the resources, search efficiently into the different fields across datasets, and run experiments on several sets at once. Metaphor datasets can be added by their authors or by anyone if their license allows it. In this section, we review the existing dataset formats and our unification choices. The website database currently contains 25 datasets. We select twelve metaphor datasets with open licenses and diverse encoded information to illustrate our unification process and as representative examples for our repository. Their characteristics are summarized in Table 1 and in more details in Appendix Table 5.
One tagged expression per example.
@@
Many datasets in NLP contain labels for binary classification or paraphrasing of one delimited expression within a given context. The expressions considered might be a single token, a multi-word expression (MWE), a phrase or a compound. The task of their original release paper may be to decide if the marked expression is used metaphorically or literally, or to find a correct literal paraphrase of the metaphoric expression, among other possible analysis and tasks. The context may consist in one single word, in the case of Adjective-Noun pairs datasets, as for example in Gutiérrez et al. (2016) or Tsvetkov et al. (2014). They often consist in one sentence such as Mohammad et al. (2016) or Cardillo et al. (2010, 2017) or Jankowiak (2020). Longer contexts are sometimes directly provided, such as in Tayyar Madabushi et al. (2022b) or Haagsma et al. (2020) who include two sentences before and after the sentence containing the marked Potential Idiomatic Expression (PIE), or indirectly provided with a reference to the original document, as for example in Joseph et al. (2023) who share references to newspaper articles or Tu and Roth (2012) who link the annotated examples to sentence pointers in the British National Corpus.
Multiple tags per example.
@@
Steen (2010) defined an annotation guideline that is widely used in metaphor studies, the MIPVU procedure, together with the release of the VU Amsterdam Corpus (VUAC), the largest existing metaphor corpus. Other examples of English open datasets created following MIPVU, that are in MetaphorShare, are Nacey et al. (2019b) and Julich-Warpakowski (2022) on the Music Criticism domain. Many other MIPVU datasets exist in various languages with adaptations of the guideline Nacey et al. (2019a), for example Nacey (2022) for Scandinavian. In this framework, every token in a document is labelled as metaphoric or literal.
@@
Different versions of the VUAC suiting NLP tasks, adapted for binary classification tasks Leong et al. (2018), or including additional ratings of novelty Parde and Nielsen (2018) have been released over the years. In the evaluation section of this paper, Section 5, we use a modified version of the VUAC corpus, adapted for binary classification (VUAC_BO) as described in Boisson et al. (2023).
@@
Beyond metaphorical/literal labels of expressions in context, metaphor datasets may contain more or less common additional information. There are scalar variables such as metaphoricity/figurativeness ratings Cardillo et al. (2010); Dunn (2014); Katz et al. (1988), confidence scores, concreteness scores, novelty scores Parde and Nielsen (2018), level of emotion scores Mohammad et al. (2016), frequency in corpora… They may also contain categorical variables such as PoS, textual source and genre, metaphor type, source and target concept/domains information Gordon et al. (2015). Figure 1 shows an example of the additional information encoded in the Mohammad et al. (2016) dataset for one instance, as diplayed on MetaphorShare. Some datasets, for example the LCC Mohler et al. (2016) and Gordon et al. (2015), also annotate the fragment of the context that carries the lexical information about the target of a metaphor (e.g. in the phrase ocean of happiness, the metaphoric expression is ocean and the target lexical cue of the metaphor is happiness).
@@
To simplify the search into various datasets and their handling for further system development, we require a specific input format for uploading a new resource on the website. MetaphorShare currently accepts CSV files having minimally one column named tagged_text. The column must contain text with at least one tagged expression per line. The accepted tag set is currently: < m: metaphoric, l: literal, t: target, a:semantically anomalous, u: free tag>. Each cell must minimally contain one tag and multiple tags are allowed, making possible the upload of MIPVU datasets without sentence duplication, e.g. I <m>swim</m> today in an <m>ocean</m> of <t>happiness</t>. Additional information may be added in freely named columns of the CSV file. Sentence indices can therefore be added to preserve the sentence order in a document and include metaphor in discourse datasets. Redundant fields across datasets such as PoS information have recommended names provided in the website guideline and we will see in Section 4.2 that the manual validation step by the website administrators allows a discussion with the user/author for further field names unification if needed. If any author of a dataset encounters issues to reformat a dataset into our expected format, we invite her or him to contact us by email for support.
@@
Author(s) Authors of the paper
Year Paper’s publication year
@@
Source Source corpora of the labelled instances
@@
Comments Additional description and comments
@@ 4 MetaphorShare: The Website
MetaphorShare general architecture is described in Section 4.1. The website is organized around four main pages: an uploading page (Section 4.2), a dataset catalog (Section 4.3), a page for searching into the datasets records (Section 4.4), and an online annotation tool (Section 4.5).
@@ 4.2 Uploading a new dataset
The required and optional fields to be filled in at uploading time are presented in Table 2. For example, if a dataset is associated with one or several publications, the references should be added in the appropriate fields, preferably in the BibTeX format. The information provided will be displayed in the dataset catalog (c.f. Figure 2 in the Appendix), and will help to further categorize the datasets in the future versions of the website.
@@
After validating the form, an immediate automatic file format check on the data is done and feedback is returned to the user for acceptance or rejection of the file. Outputs of this step can be seen in Appendix Figure 3 & 4. In a second manual validation step, the website administrators check the license information, dataset reference, and eventually suggest modifications of CSV free columns names, that contain the dataset specific additional information, in an email interaction with the user, for database field unification purposes.
@@
Once a dataset passes the two validation steps, it is stored in the main repository database. This ElasticSearch database has two indices: the Datasets index similar to Table 2 fields stores all the user provided metadata, and information extracted during the processing of the data file such as the dataset size and the label distribution. The second one, the Potentially Metaphoric Expressions, PME index, contains the output of the data file parse, as shown Table 3. Each textual labelled example is linked to the Dataset index.
@@ 4.3 Dataset Catalog
The main purpose of this page is to display the main features of each dataset included in the repository. The fields filled at uploading time, including the license chosen, are displayed in the dataset catalog page. Statistics computed from the input file are also shown and plotted, in particular the number of instances, label set and distribution. A partial view of a dataset summary in the catalog is shown in the Appendix, Figure 2.
@@ 4.4 Search Page
The records can be accessed and downloaded through a search page, as shown in Figure 1. The search fields include optional selections of datasets and languages from drop-down lists, restricted search for a label (e.g. metaphorical or literal), and keyword search into the PME text (with exact and fuzzy match), or into the entire text. One phrase or sentence is shown by default on the result page, but longer context can be accessed in one click, as well as all the additional information attached to a record. [^2] The organization of the catalog and the fine-grained search options are likely to evolve in future versions of MetaphorShare. The result of any search can be downloaded as a CSV file.
4.5 Online data annotation
On the Edit/Label Dataset page, one can choose to directly input some text into the interface of the annotation tool, or alternatively to upload a table (in a CSV or Excel format) containing unannotated or partially annotated text. An input can then be labelled from scratch or edit labels by pairing highlighted expressions with tags (c.f. Figure 5 in the Appendix). While the MetaphorShare tagset is predefined, custom tags can also be created. The resulting labelled data (c.f. Appendix Figure 6) can then be downloaded as a CSV, or directly uploaded in MetaphorShare following the procedure described in Section 4.2. This online annotation functionality facilitates the creation of resources immediately compatible with our accepted formats, and speeds up the correction of CSV files that do not pass the automatic validation step of the dataset upload procedure. It is also an accessible tool for annotators without any computer science background.
@@ 5 Evaluation
As a sanity check, we first verified that MetaphorShare uploading, database insertion and search functionalities are working properly with the datasets presented in Table 2.
Then, as a case study we perform a cross-dataset metaphor identification analysis to illustrate a possible usage of MetaphorShare for NLP research. Given an input expression, the task consists of deciding whether the expression is a metaphor or not. Different projects label metaphors for different purposes. MetaphorShare makes it easy to fine-tune models on relevant dataset to further support project-specific automatic or semi-automatic labeling.
@@
In order to maintain a training set of the same size for all datasets, we randomly sampled 800 examples from each set to create the training sets. Because our datasets from psycholinguisics are small and similar in their creation method, we grouped them into one single set called J&C in this experiment. RoBERTa base models Liu et al. (2019) are then finetuned independently on the 10 datasets and tested on each of them. Similarly to the experiments in Boisson et al. (2023), for hyperparameter optimisation, we rely on the Bayesian Optimization with the Hyperband (BOHB) algorithm Falkner et al. (2018) available in RayTune Liaw et al. (2018), with an identical hyperparameter search space and 25 trials.